OpenStudy (anonymous):

is the function f(x)=3x^4+4x+8 even odd or niether. show proof

OpenStudy (amistre64):

if f(-x) = f(x); its even if f(-x) = -f(x) it is odd if it aint those, then its neither

OpenStudy (amistre64):

f(x)=3x^4+4x+8 f(-x)=3(-x)^4+4(-x)+8 = 3x^4 - 4x + 8 f(-x) NOT= f(x); so it aint even

OpenStudy (amistre64):

f(x)=3x^4 +4x +8 -f(x) = -3x^4 -4x -8 f(-x) NOT= -f(x) ; it aint odd its got to be neither

OpenStudy (anonymous):

thank u

OpenStudy (amistre64):

yep :)